OpenStudy (sleepyjess):

Simonne has been following the stock ratings of Carter's Pork Bellies. She has noticed that the pattern can be modeled by a cosine function. During her observations, she has found the stock to have a maximum of x = 6, minimum of x = −2, and period of π. Which of the following cosine functions could model Simonne's stock rating?

OpenStudy (sleepyjess):

f(x) = 4 cos 2x − 4 f(x) = 6 cos pi over 2x + 2 f(x) = 6 cos pi over 2x − 4 f(x) = 4 cos 2x + 2

OpenStudy (sleepyjess):

@phi

OpenStudy (phi):

The first step could be, what is the range of the values from low to high ?

OpenStudy (sleepyjess):

6?

OpenStudy (phi):

x = 6, minimum of x = −2 so the distance from min to max is ?

OpenStudy (sleepyjess):

8

OpenStudy (phi):

a cosine has a min of -1 and a max of +1, and a range of 2 your cosine range is 8. that means multiply the cosine by 4 4 cos(stuff) will range from -4 to +4 or a total of 8

OpenStudy (phi):

however, we have to tweak that, so it ranges from -2 to +6 I would add +2 to it, like this: 4 cos(stuff) + 2 now it will range from -2 to +6

OpenStudy (sleepyjess):

ok thank you so much for your help! :-)

OpenStudy (phi):

It looks like we did not need to use the period to get to the correct answer. But if we had to, we would do this: inside a cosine, (the "stuff" ), write it as \[ \frac{2 \pi}{T} x \] where T is the period they tell use T is π so the inside should be \[ \frac{2 \pi}{\pi} x = 2x \] the full answer is \[ 4 \cos(2x)+2 \]